Daily Report

CPUC/ANF Environmental Monitoring Activities

Pacific Pipeline Construction Project

Aspen Environmental Group Monday, January 18, 1999

The following is a brief summary of the environmental monitoring activities that were completed between January 18, 1999 1:00 and 1:00 on January 19, 1999. This summary is based on reports directly received from the field and has not been edited. If you have any questions please contact (818) 597-3407. Please note the item number and the correction that should be made. This report is intended to serve as a communication document and as an instrument for quality control.

Construction Highlights

01/18/99 (Monday)

Segment 1 (Emidio Pump Station to Copco Road): Report for 18 January 1999: Restoration efforts began along Rock Road today. Testing crews ran pigs between the Cuddy Creek Tributary test site and Grapevine Station, as well as from Grapevine Station to Emidio Pump Station. Crews worked on the tie in at the south side of Grapevine Pump Station late in the day. Welding on pups and surplus pipe continued along Highway 166. Coating crews finished work on the tie in at Copco Road. General clean up was done segment wide. Backfill, compaction, and clean up progressed on upper Grapevine Road, near the bore site.

Segment 5 (Lynwood to Long Beach): Report for 1-18-99. Crews cut out a short segment of the 20'' pipe and began welding the 16'' tie in for the Edison section. Further, water from the hydro test was removed after completion of testing and hauled to a local water plant. In addition, crews removed soil from the receiving pipe and set up a small scaffold to support the crews. Welding continued south of the Tosco refinery on the Huntway section with crews connecting 40' lengths of pipe. Meanwhile, excavation of the bore pit within the refinery continued while crews removed soil from previously buried pipe. This was required after x-ray films were misplaced for several welds. In addition, crews progressed with the meter electrical power, pouring red concrete around electrical conduit and general clean up. Segment 4B (Lynwood to El Segundo): Crews installed headers for the Lynwood station roof, washed walls and removed debris from the right of way.

Inspection Summar

In the past 24 hours, 2 inspection event(s) were reported for 2 spread(s). Of these inspections, 0 were non-compliance event(s). Also 0 non-compliance resolution report(s) was resolved.

Non-Compliance Reports (NCR's) Issued

No new non-compliance reports were issued during the last reporting period.

Non-Compliance Resolution Reports (NCRR's) Issued

No new non-compliance reports were issued during the last reporting period.

Summary of Follow-up Items for Today

No follow-up items were reported for today.

Back to Daily Reports

Back to MMCRP Home Page